Find the value of M. M -:-26=3 1/4. I tried multiplying 3 1/4 by 26 but i couldent find a way how.
Answer by ilana(307)

If you know how to multiply fractions and whole numbers, make 3 1/4 into an improper fraction. 3*4=12 and 12+1=13, so 3 1/4 = 13/4.
Now multiply that by 26 just like you were.